OpenStudy (heavymetal356):

Plz someone help!!! A rock is thrown vertically upward from the ground, with an initial velocity of 64 ft/sec. Its position function is s(t) = –16t^2 + 64t. What is its velocity in ft/sec when it hits the ground? 0 –64 –32 –16

OpenStudy (heavymetal356):

plz help me I am really stuck would the answer be 0?

OpenStudy (slipknot306):

A. i think

OpenStudy (heavymetal356):

thx @slipknot how about this one? Use the graph of f "(x) below to state x-coordinates of the inflection points for the graph of f(x). graph is an inverted parabola with x intercepts at x equals 0 and x equals 4 and vertext at x equals 2. 0 only 4 only 0 and 4 2 only

OpenStudy (heavymetal356):

A water tank has the shape of an inverted right circular cone with base radius 3 meters and height 6 meters. Water is being pumped into the tank at the rate of 12 meters3/sec. Find the rate, in meters/sec, at which the water level is rising when the water is 2 meters deep. Give 2 decimal places for your answer. Type your answer in the space below. If your answer is a number less than 1, place a leading "0" before the decimal point (ex: 0.35).

OpenStudy (slipknot306):

C.

OpenStudy (slipknot306):

for the next one

OpenStudy (heavymetal356):

I am not really sure for the next one, could you walk me through it?

OpenStudy (slipknot306):

lol i dont know either x'd give meh a sec

OpenStudy (slipknot306):

let the radius of the water level be r m let the height of the water be h m make a sketch and by similarity: h/r = 6/3 h/r = 2 h = 2r or r = h/2 V = (1/3)π r^2 h = (1/3)π (h^2/4)(h) = (1/12)π h^3 dV/dt = (1/4) h^2 dh/dt so when h= 2 and dV/dt = 12 12 = (1/4)((4)dh/dt dh/dt = 12 m/sec

OpenStudy (slipknot306):

there u go

OpenStudy (heavymetal356):

i'm sorry but I can't use that. It's not right.

OpenStudy (slipknot306):

x'd well its wut google said

OpenStudy (heavymetal356):

do you know the correct answer?

OpenStudy (slipknot306):

um...no x'd

OpenStudy (heavymetal356):

well, I submit my test and got a 77%, I missed 5.